OpenStudy (anonymous):

Calculate f (6) a) f (x) = 7x - 20 b) f (x) = 2x ^ 2 c) f (x) = 11500 * 1.035 ^ x

OpenStudy (anonymous):

just put 6 in place of x

OpenStudy (nirmalnema):

a) 7(6)-20 =42-20 =22 b) 2(2)^2 2*4=8 c) 11500 * 1.035 ^ 2 =12319.08

OpenStudy (nirmalnema):

@sarapetrini do u understand